NandishSS wrote:Three cannons are firing at a target. If their individual probabilities to hit the target are 0.3, 0.4, and 0.5 respectively, what is the probability that none of the cannons will hit the target after one round of fire?
A. 0.06
B. 0.12
C. 0.21
D. 0.29
E. 0.94
[spoiler]OA: C[/spoiler]
HI Experts
Why is below ans incorrect?
P [None] = 1 - P [all hit target]
= 1 - (0.3) (0.4) (0.5)
= 0.94
Thanks
Nandish
P[all shots hit target] is not the complement of P[no shots hit target]
If what is the opposite of "none of the shots hit the target"?
Well, it's can be 1 shot hits target or 2 shots hit target or all 3 shots hit target.
So, if we're going to use the complement here, we must consider all 3 cases.
That is:
P(no shots hit target) = 1 - P(1 shot hits target OR 2 shots hit target OR all 3 shots hit target)
Lots of work!!!
Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
